A particle describes simple ahrmonic motion with amplitude of 5 cm and period of 0.2 sec. Calculate the velocity and acceleration of the particle when the displacement is 0 cm


1
Expert's answer
2021-12-10T13:21:10-0500

For simple harmonic motion


"x(t)=A\sin(\omega t)\to x(t)=0.05\cdot\sin(10\pi t)"


"v(t)=0.05\cdot(10\pi)\cdot\cos(10\pi t)=0.5\pi\cdot\cos(10\pi t)"


"a(t)=-0.5\pi\cdot10\pi\cdot\sin(10\pi t)=-5\pi^2\cdot\sin(10\pi t)"


If "x(0)=0" then "t=0" . So,



"v=0.5\cdot3.14=1.57\ (m/s)" ;


"a=0" .